OpenStudy (anonymous):

Find dy/dx if y = sin (x+y)? Here's my work, but I think it's wrong... d/dx [y] = d/dx [sin (x+y)] dy/dx = cos (x+y) * d/dx (x+y) Please help...

OpenStudy (math&ing001):

Nothing wrong in your answer, you just didn't finish yet. Find d/dx (x+y).

OpenStudy (anonymous):

dy/dx (x+y) = (1+1) = 2? So... would it be dy/dx = 2 cos (x+y)? wait... that's not one of the answer choices :'( Help me out here, please?

OpenStudy (math&ing001):

d/dx (x+y) = 1

OpenStudy (anonymous):

\[\frac{ d }{ dx } (x + y) = 1 + \frac{ dy }{ dx }\]

OpenStudy (math&ing001):

Sorry my bad ! @niksva is right.

OpenStudy (anonymous):

Thank you guys! So dy/dx = cos (x+y) * 1+dy/dx... and do I move 1+dy/dx to the left?

OpenStudy (math&ing001):

You distribute first dy/dx=cos(x+y)+cos(x+y)*dy/dx. Then yeah you move cos(x+y)*dy/dx. to the left.

OpenStudy (anonymous):

I think I got it...! Is the answer dy/dx = cos (x+y) / 1-cos (x+y)??

OpenStudy (math&ing001):

That's right !

OpenStudy (anonymous):

Ah, finally! Thanks so much!!!! :) :) To niksva too! :)

OpenStudy (math&ing001):

Your welcome

OpenStudy (anonymous):

pleasure :-P